# Omni-Think: Scaling Cross-Domain Generalization in LLMs via Multi-Task RL with Hybrid Rewards

### v1

@@ Abstract
The advancement of general-purpose artificial intelligence relies on large language models (LLMs) that excel across a wide range of tasks, from structured reasoning to creative generation. However, post-training methods like Supervised Fine-Tuning (SFT) often struggle with generalization, favoring memorization over transferable learning. In this work, we introduce Omni-Think, a unified reinforcement learning (RL) framework that enhances LLM performance across diverse tasks by combining rule-based verifiable rewards with generative preference signals via LLM-as-a-Judge evaluations. Our approach enables consistent optimization across task types and scales RL-based training to subjective domains. We further investigate training strategies, demonstrating that a curriculum-based progression that orders tasks from structured to open-ended improves performance and reduces forgetting. Experimental results across four domains reveal that curriculum learning improves performance by 5.2% over joint training and 9.1% over model merging. These results highlight the importance of task-aware sampling and hybrid supervision in scaling RL-based post-training for general-purpose LLMs.
@@ 1 Introduction
Figure 1: OMNI-THINK Framework Overview. A unified multi-task RL framework for training a single LLM across structured and open-ended tasks. (1) Backward Transfer quantifies task forgettability and informs (2) a curriculum that schedules tasks from least to most forgettable. (3) Sampled tasks construct prompts. (4) The LLM generates completions, which are (5) scored by hybrid rewards: verifiable (e.g., test case execution), generative (e.g., LLM-as-a-Judge), and auxiliary (e.g., format adherence), among others. (6) The policy is updated via Multi-Task GRPO. The process repeats via iterative updates (bottom loop).
@@
To address this, we propose Omni-Think, a unified RL framework that enables LLMs to learn from both rule-based and generative supervision across tasks. Building on Reinforcement Learning with Verified Reward (RLVR), our method integrates symbolic verifiers with LLM-as-a-Judge evaluations (Zheng et al., 2023; Zhang et al., 2025) to handle subjective tasks. We further show that curriculum training, which sequences tasks from structured to open-ended based on Backward Transfer (BWT) measurements of task forgettability, significantly improves generalization and reduces forgetting. Across four domains, it achieves average gains of 5.2% and 9.1% over joint training and model merging, respectively.
Our key contributions are: (1) We propose Omni-Think, a unified framework that supports both verifiable and generative supervision under a single policy, scaling RL across four diverse domains. (2) We show that LLM-as-a-Judge enables scalable reward signals for open-ended tasks, extending GRPO beyond rule-based domains. (3) We demonstrate that not all multi-task training strategies are equally effective. Curriculum-based task scheduling outperforms uniform sampling and improves generalization in multi-task RL.
@@ 4 Methodology
4.1 Omni-Think Framework Overview
We introduce Omni-Think, a unified post-training framework that aligns LLMs to a wide range of tasks via reinforcement learning, extending RLVR beyond structured domains like math and code to include general and open-ended tasks.
@@ 4.1 Omni-Think Framework Overview
For concise generative tasks with known answers (e.g., General question answering (QA)), we transform the questions into open-ended prompts and embed the distractor responses, which are plausible but incorrect candidate answers, within the context. Rather than labeling options, we instruct the model to reason with <think></think> and produce answers in <answer></answer> format. A response is assigned a binary reward based on whether it matches any reference answer via string comparison or set membership, promoting semantic grounding and discouraging pattern memorization.
@@
For subjective tasks lacking ground truth (e.g., dialogue, writing), we use an LLM-as-a-Judge  (Chen et al., 2025) to assign a scalar reward based on rubric-aligned pairwise preferences between candidate outputs. This enables learning in domains where symbolic correctness is insufficient or intractable.
@@ 5.1 Training Datasets
We curate a multi-domain training dataset covering Math, Coding, General QA, and Creative Writing, with each domain selected to support hybrid reward functions and robust evaluation. For Math, we begin with the OpenR1-Math (HuggingFace, 2025) dataset, retaining only word problems and excluding questions that require visual reasoning. We further subsample 12,000 examples to fit our compute budget. For Coding, data is sourced from the code-r1-12k (Liu & Zhang, 2025) dataset, with outliers exceeding 1024 tokens removed. Each entry includes a code prompt and JSON-formatted unit tests for automatic validation. For General QA, we extract 5,500 examples from SuperGPQA (M-A-P, 2025), downsampled proportionally by question category. Each sample comprises a factual question paired with a plain-text answer. The Creative Writing domain leverages 6,650 conversations from Nitral AI’s ShareGPT dataset (Nitral‑AI, 2024), focused on one-turn completions. Samples exceeding two dialogue turns are filtered out, and responses are judged via an LLM-as-a-Judge framework.
@@ 6.1 Main Results: Scaling Multi-Task LLM Post-Training with Omni-Think
We evaluate Omni-Think across four diverse domains: Coding, Math, General QA, and Creative Writing, to assess how reinforcement learning with rule-based verifiable rewards and generative supervision supports multi-task generalization. Figure 2 shows that Curriculum-Guided Multi-Task GRPO consistently yields the best results. Table 2 further details how these gains vary by benchmarks.
@@
These results support our central hypothesis: The Omni-Think Training Framework, with the help of Curriculum MT, enables a single unified policy to scale across structured and open-ended tasks alike, without relying on interleaving RLVR on reasoning tasks and fine-tuning non-reasoning tasks.
@@ 7 Conclusion
We presented OMNI-THINK, a unified framework that extends LLMs to handle both structured and open-ended tasks within a single policy. By integrating reinforcement learning with both rule-based verifiable rewards and generative preference-based supervision, our method improves generalization while addressing challenges such as forgetting and task interference. Our findings show that effective multi-task LLM post-training depends not only on reward design but also on how tasks are sequenced and optimized together. By ordering tasks from structured to open-ended domains based on backward transfer, OMNI-THINK minimizes forgetting and achieves consistent gains across diverse reasoning and generative tasks. Overall, it takes a step toward more general-purpose LLMs that can learn effectively from both verifiable and subjective feedback, bridging structured reasoning, open-ended question answering, and creative generation within a single, unified post-training framework.
@@ A.1 Reward Estimation
Omni-Think employs a hybrid reward system combineing rule-based correctness (math, code, QA) with preference-based supervision (creative writing) in a unified RL framework. We define task-specific reward functions as $r_{k}(o,\phi_{k})$, where $o$ denotes the model output and $\phi_{k}$ represents auxiliary task-specific inputs required for evaluation—such as reference answers, unit tests, or candidate responses for comparison. Each reward function captures domain-relevant correctness criteria, assessing whether $o$ satisfies symbolic constraints, passes execution tests, or is preferred over alternatives under subjective evaluation. While some rewards (e.g., math and code) are strictly deterministic, others, such as LLM-as-a-Judge comparisons, are inherently stochastic but executed at low decoding temperature to ensure stable and consistent supervision. All reward functions are designed to be domain-aware and automatable, supporting scalable reinforcement learning across both structured and generative tasks.

### v2

# Omni-Thinker: Scaling Cross-Domain Generalization in LLMs via Multi-Task RL with Hybrid Rewards

@@ Abstract
The advancement of general-purpose artificial intelligence relies on large language models (LLMs) that excel across a wide range of tasks, from structured reasoning to creative generation. However, post-training methods like Supervised Fine-Tuning (SFT) often struggle with generalization, favoring memorization over transferable learning. In this work, we introduce Omni-Thinker, a unified reinforcement learning (RL) framework that enhances LLM performance across diverse tasks by combining rule-based verifiable rewards with generative preference signals via LLM-as-a-Judge evaluations. Our approach enables consistent optimization across task types and scales RL-based training to subjective domains. We further investigate training strategies, demonstrating that a curriculum-based progression that orders tasks from structured to open-ended improves performance and reduces forgetting. Experimental results across four domains reveal that curriculum learning improves performance by 5.2% over joint training and 9.1% over model merging. These results highlight the importance of task-aware sampling and hybrid supervision in scaling RL-based post-training for general-purpose LLMs.
@@ 1 Introduction
Figure 1: OMNI-THINKER Framework Overview. A unified multi-task RL framework for training a single LLM across structured and open-ended tasks. (1) Backward Transfer quantifies task forgettability and informs (2) a curriculum that schedules tasks from least to most forgettable. (3) Sampled tasks construct prompts. (4) The LLM generates completions, which are (5) scored by hybrid rewards: verifiable (e.g., test case execution), generative (e.g., LLM-as-a-Judge), and auxiliary (e.g., format adherence), among others. (6) The policy is updated via Multi-Task GRPO. The process repeats via iterative updates (bottom loop).
@@
To address this, we propose Omni-Thinker, a unified RL framework that enables LLMs to learn from both rule-based and generative supervision across tasks. Building on Reinforcement Learning with Verified Reward (RLVR), our method integrates symbolic verifiers with LLM-as-a-Judge evaluations (Zheng et al., 2023; Zhang et al., 2025) to handle subjective tasks. We further show that curriculum training, which sequences tasks from structured to open-ended based on Backward Transfer (BWT) measurements of task forgettability, significantly improves generalization and reduces forgetting. Across four domains, it achieves average gains of 5.2% and 9.1% over joint training and model merging, respectively.
Our key contributions are: (1) We propose Omni-Thinker, a unified framework that supports both verifiable and generative supervision under a single policy, scaling RL across four diverse domains. (2) We show that LLM-as-a-Judge enables scalable reward signals for open-ended tasks, extending GRPO beyond rule-based domains. (3) We demonstrate that not all multi-task training strategies are equally effective. Curriculum-based task scheduling outperforms uniform sampling and improves generalization in multi-task RL.
@@ 4 Methodology
4.1 Omni-Thinker Framework Overview
We introduce Omni-Thinker, a unified post-training framework that aligns LLMs to a wide range of tasks via reinforcement learning, extending RLVR beyond structured domains like math and code to include general and open-ended tasks.
@@ 4.1 Omni-Thinker Framework Overview
For language tasks with known or extractable ground-truth answers such as general question answering (QA), we reformulate queries into open-ended prompts and incorporate distractor responses (LLM-generated plausible but incorrect answers) into the context. Instead of labeling options, we prompt the model to reason using the <think></think> format and to output answers within <answer></answer> tags. Responses are evaluated with a binary reward based on string matching or set membership against reference answers, thereby encouraging semantic grounding and mitigating shallow pattern memorization. We find that conditioning the LLM on a diverse set of candidate options, including one correct answer and multiple distractors, is key to steadily improving general-domain reasoning while reducing susceptibility to random guessing or reward hacking, compared to directly prompting the model to generate open-ended answers during training without the augmented context.
@@
For subjective tasks lacking ground truth (e.g., dialogue, writing), we use an LLM-as-a-Judge  (Chen et al., 2025) to assign a scalar reward based on rubric-aligned pairwise preferences between candidate outputs. This enables learning in domains where symbolic correctness is insufficient or intractable. This prompt-based approach leverages recent advances in the general reasoning capabilities of LLMs, using generated chain-of-thoughts to elicit a ternary reward signal, preferred, tie, or dispreferred, without requiring large-scale preference data collection and reward model training.
@@ 5.1 Training Datasets
We curate a multi-domain training dataset covering Math, Coding, General QA, and Creative Writing, with each domain selected to support hybrid reward functions and robust evaluation. For Math, we begin with the OpenR1-Math (HuggingFace, 2025) dataset, retaining only word problems and excluding questions that require visual reasoning. We further subsample 12,000 examples to fit our compute budget. For Coding, data is sourced from the code-r1-12k (Liu & Zhang, 2025) dataset, with outliers exceeding 1024 tokens removed. Each entry includes a code prompt and JSON-formatted unit tests for automatic validation. For General QA, inspired by SuperGPQA (M-A-P, 2025), we subsample 5,500 queries from the benchmark dataset, proportionally by question category. Each sample comprises a factual question paired with a plain-text answer. We then generate 15 additional confusion options while making sure the uniqueness of correctness by prompting an LLM. The Creative Writing domain leverages 6,650 conversations from Nitral AI’s ShareGPT dataset (Nitral‑AI, 2024), focused on single-turn completions. Samples exceeding two dialogue turns are filtered out, and responses are judged via an LLM-as-a-Judge framework.
@@ 6.1 Main Results: Scaling Multi-Task LLM Post-Training with Omni-Think
We evaluate Omni-Thinker across four diverse domains: Coding, Math, General QA, and Creative Writing, to assess how reinforcement learning with rule-based verifiable rewards and generative supervision supports multi-task generalization. Figure 2 shows that Curriculum-Guided Multi-Task GRPO consistently yields the best results. Table 2 further details how these gains vary by benchmarks.
@@
These results support our central hypothesis: The Omni-Thinker Training Framework, with the help of Curriculum MT, enables a single unified policy to scale across structured and open-ended tasks alike, without relying on interleaving RLVR on reasoning tasks and fine-tuning non-reasoning tasks.
@@ 7 Conclusion
We presented Omni-Thinker, a unified reinforcement learning framework that enables large language models to handle both structured and open-ended tasks under a single policy. By combining rule-based verifiable rewards and generative preference-based supervision, our method improves generalization while mitigating forgetting and interference. Our findings show that effective multi-task LLM post-training depends not only on reward design but also on how tasks are sequenced and optimized together. Ordering tasks from structured to open-ended domains based on backward transfer reduces forgetting and enhances cross-domain performance. Overall, Omni-Thinker advances the goal of general-purpose LLMs that can learn from both verifiable and subjective feedback, bridging structured reasoning, open-ended question answering, and creative generation in a single post-training framework.
@@ A.1 Reward Estimation
Omni-Thinker employs a hybrid reward system combineing rule-based correctness (math, code, QA) with preference-based supervision (creative writing) in a unified RL framework. We define task-specific reward functions as $r_{k}(o,\phi_{k})$, where $o$ denotes the model output and $\phi_{k}$ represents auxiliary task-specific inputs required for evaluation—such as reference answers, unit tests, or candidate responses for comparison. Each reward function captures domain-relevant correctness criteria, assessing whether $o$ satisfies symbolic constraints, passes execution tests, or is preferred over alternatives under subjective evaluation. While some rewards (e.g., math and code) are strictly deterministic, others, such as LLM-as-a-Judge comparisons, are inherently stochastic but executed at low decoding temperature to ensure stable and consistent supervision. All reward functions are designed to be domain-aware and automatable, supporting scalable reinforcement learning across both structured and generative tasks.